OpenStudy (anonymous):

how do i find the limit as n approaches infinity of (1/4 + 1/2n + 1/4n^2)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Well, as n becomes arbitrarily large 1/2n and 1/4n^2 both tend to zero. So your answer is just 1/4 since it is not dependent on n.

OpenStudy (anonymous):

it us 1/4 the next two terms reduce to zero.

OpenStudy (amistre64):

unless.......... maybe........... nah, thats right.

OpenStudy (amistre64):

big bottomed frations are very very skinny

OpenStudy (anonymous):

amistre questioning me?!?!? D:< J/k j/k <33 lol

OpenStudy (amistre64):

1 ----------- = .000...00001 10000...000

OpenStudy (amistre64):

youre icon reminds me of one of my personalities ;)

OpenStudy (mathteacher1729):

1) http://www.wolframalpha.com 2) type " limit as n approaches infinity of (1/4 + 1/2n + 1/4n^2)" 3) Click "show steps" Truly, we live in the future.

OpenStudy (amistre64):

im old, new things scare me ..... now wheres my sliderule?

OpenStudy (anonymous):

mathteacher no one wants to know how to plug everything in a calculator. Some people want to actually understand what they're doing. So take your advertisement somewhere else :)

OpenStudy (anonymous):

show steps? ever thing goes to zero except the number.

OpenStudy (anonymous):

Wolfram is a LAST resort if you're stuck.

OpenStudy (anonymous):

actually i like it for graphing. but to answer this question use your mental wolfram

OpenStudy (amistre64):

wolfram is cool for alot of stuff; but mines stuck on multiplication flip cards

OpenStudy (amistre64):

i know how to get to the: 1: quit 2: goto screen on my ti-83

OpenStudy (anonymous):

like asking what \[lim_{x->\infty} \frac{x^2+2x}{x-1}\] by a graph!

OpenStudy (mathteacher1729):

Agreed Satellite. I like "use your mental wolfram". The explanations it provides are not always straightforward or the most efficient. It gets better all the time, but nothing is a substitute for developing a good numerical / algebraic intuition and understanding. :)

OpenStudy (mathteacher1729):

http://www.math.psu.edu/ug/courses/math140 has GREAT limits review which helps develop this intuition by comparing relative rates of growth for different kinds of functions. I would call it a MUST READ. :)

OpenStudy (anonymous):

how about straight up visualization especially for a problem like this. every blasted thing goes to zero except the number.

OpenStudy (mathteacher1729):

Whups, they changed the link. Here it is: http://www.math.psu.edu/files/141rates1.pdf That's the relative rates of growth I was referring to.

OpenStudy (mathteacher1729):

To visualize this, type "plot (1/4 + 1/2n + 1/4n^2) from n = 1 to n = 100" in wolframalpha